OpenStudy (anonymous):

how many grams of O2 gas would be requied to react completely with 4.86g of Mg. balance equation is: 2Mg+O2-->2MgO

thomaster (thomaster):

As you see 2 mol magnesium reacts with 1 mol oxygen to form 2 mol magnesiumoxide. What you need do to is convert grams of Mg to mol Mg. Magnesium has a molecular weight of 24.305 g/mol To calculate how many mol 4.86 g of magnesium is, you divide 4.860 by 24.305\[\frac{ 4.860 }{ 24.305 }=0.199 mol\]The ratio is 2:1:2, but we only need the first two. 2 mol Mg is used to react with 1 mol O2. So when you have 0.199 mol Mg, it will react with 0.199/2= 0.099 mol O2 Last thing to do is convert 0.099 mol O2 to grams. O2 has a molecular weight of 31.998 g/mol To calculate how many grams 0.099 mol oxygen is you multipy it by the molecular weight. So 0.099*31.998= 3.168 gram 4.860 grams of Mg will react with 3.168 grams of O2

thomaster (thomaster):

Also you can find molecular weights on a periodic table http://www.ptable.com is a nice one

OpenStudy (anonymous):

First you must find mole of Mg mole of Mg = mass/Ar = 4.86/24.3 = 0.2 Mole of oxygen = 1/2 x 0.2 = 0.1 mass of oxygen = 0.1 x 32 = 3.2 gram